OpenStudy (anonymous):

Eliminate the parameter. x = t - 3, y = t2 + 5

OpenStudy (anonymous):

solve for t using one of the equations t = -x -3 Plug it in the other one y = (-x-3)^2 + 5 Simplify and get your answer

OpenStudy (anonymous):

x= t-3 to rid -3 form right side so we add 3 to both sides so we get x +3 = t - 3 + 3 x + 3= t + 0 x+3 = t now plug this value for t = x + 3 in the y = t^2 +5 so we get y = (x+3)^2 + 5 this is the final answer.

OpenStudy (anonymous):

@keti and @alias Thank you so much!!!♥♥

OpenStudy (anonymous):

@keti has the right answer :P I made a mistake when isolating t in the first eqn.

OpenStudy (anonymous):

okay. do you have any more questions